OpenStudy (anonymous):

What is the volume of a sphere with a surface area of 25π yd2?

OpenStudy (anonymous):

25 yards2 ???? If so...., Area sphere = 4*pi*r2 You have area, so...., 25 yrd2 = 4*pi*r2 r = sqrt(25/4*pi) = 1.41 yards in radius ----------------------------- Volume sphere = 4/3*pi*r3 Volume = 4/3*pi(1.41 yrd)3 = 12 cubic yards